Tohban Report 2015-05-13

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 6 May 2015
End Date: 13 May 2015
Tohban: Juan C. Martinez Oliveros
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

Solar activity was low this week. There are currently 12 ARs on the disk with the probability of some M flares in the following days. Active regions 12339 was the most active during this week.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    49     1     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    22     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0    10     0     0

There were RHESSI flares/GOES flares 285 / 50 over the time range 06-May-15 13-May-15

Memory Management

The SSR fill level reached a maximum of 32% and a minimum of about 3% during the week.

Spacecraft Status

The cold plate 1 and 2 are at about 132.5 K and 131.3 K respectively.

Data Gaps

No data gaps this week

Detector issues

8 May 2015 22:33:30 UTC: The HV on detector 8 was decreased by a total of 24 counts, we went from 4020 V to 3554 V. As a result of this change the rear livetime went up, however the rear fast rates and resets increased.

15-128-22:27:41 /ihvdac detector=8, voltage=198
15-128-22:32:16 /ihvdac detector=8, voltage=190
15-128-22:33:30 /ihvdac detector=8, voltage=182 (3554 V)

12 May 2015 21:03:52 UTC: A drastic decrease in detector 8’s HV by 50 steps to about 2573 V was performed in an attempt to improve its live time.

2015-132-21:03:52 /ihvdac detector=8, voltage=132 (2573 V)

Other notes

Spacecraft Management

A spin up was performed on May 8, 2015 20:56 and 22:42 UTC.

As discussed in the previous meeting, a cryocooler rebalancing that was made on 12 May 2015. The balancer amplitude and the phase were adjusted both up and down and settled on new settings with a reduction of about 1.5 mG in acceleration readings.

The cryocooler commanding took place during the Wallops pass at 132-17:52 and again the following orbit at 132-19:30 UT. We ended up by reducing the amplitude setting by 6 counts (to 192) and increasing the phase by 2 counts (to 226), the power setting was not changed.

Since 10 Sunday 2015 the thin attenuator was stuck in (130-17:07 and 132-22:22 UT), to unlock it the fast front threshold of the 1F and 7F detectors were change to 0x40 on 12 May 2015. The attenuator came out about an hour after the threshold changes were made.

2015-132-21:01:12 /IDPLOAD VALUE=0x30  ; was 0x22

2015-132-21:01:53 /IDPLOAD VALUE=0x30  ; was 0x22

2015-132-21:03:20 /IDPLOAD VALUE=0x40  ; was 0x30

2015-132-21:10:35 /IDPLOAD VALUE=0x40  ; was 0x30

Decimation Normal/Vigorous
Night time data (fronts) No nighttime events
Night time data (rears) Taking nighttime events
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Personal tools